# On the handling of method failure in comparison studies

### v1

@@ Abstract
Comparison studies in methodological research are intended to compare methods in an evidence-based manner, offering guidance to data analysts to select a suitable method for their application. To provide trustworthy evidence, they must be carefully designed, implemented, and reported, especially given the many decisions made in planning and running. A common challenge in comparison studies is to handle the “failure” of one or more methods to produce a result for some (real or simulated) data sets, such that their performances cannot be measured in those instances. Despite an increasing emphasis on this topic in recent literature (focusing on non-convergence as a common manifestation), there is little guidance on proper handling and interpretation, and reporting of the chosen approach is often neglected. This paper aims to fill this gap and provides practical guidance for handling method failure in comparison studies. In particular, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases. We also discuss how method failure in published comparison studies—in various contexts from classical statistics and predictive modeling—may manifest differently, but is often caused by a complex interplay of several aspects. Building on this, we provide recommendations derived from realistic considerations on suitable fallbacks when encountering method failure, hence avoiding the need for discarding data sets or imputation. Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two illustrative comparison studies.
@@ 1 Introduction
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure”. Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. This is illustrated by our review of published comparison studies from classical statistics and predictive modeling in Section 2 which shows that (reported) handlings often differ greatly between comparison studies. Nevertheless, these studies are positive examples of the reporting and the handling of method failure as this issue often goes unreported in many other comparison studies. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al., , 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) reported convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility even when not reported in the manuscript. This is often clear through the use of R functions try() or tryCatch() for the handling of errors, for example.
@@
To illustrate the difficulty of handling method failure adequately, imagine conducting a comparison study in statistics or predictive modeling to compare the performance of a selection of methods. In the context of classical statistics, you likely focus on comparing the methods’ performances in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s)” and base your analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, you are likely to compare the methods’ predictive performances, that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmarking data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies”, respectively (though admitting that there may be intermediate situations, e.g. simulation studies investigating predictive performance on artificial data).
@@
During your experiments, you encounter the failure of a method in one or more of the following ways: (i) you receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) your system crashes, or (iii) the experiment runs for an excessively long time, forcing you to abort.
Table 1: Overview table of (a) a fictive simulation study and (b) a fictive benchmark study encountering method failure. This results in undefined estimates values for the simulation study (a) and undefined performance values for the benchmark study (b), marked in blue. In both cases, this complicates the aggregation of performance of the given method.
(a) Fictive simulation study: overview of estimates produced by three methods under investigation (columns) across $4$ repetitions, i.e. simulated data sets (rows), in one simulation scenario, where the true value is equal to $4$. An estimates value is obtained for each of the three methods and four simulation repetitions. Bias is commonly calculated as the mean deviation of the estimates from the true value across the $4$ repetitions. Method failure first leads to an undefined estimates value in the given repetition, which then complicates the estimation of performance of the affected method.
@@
(b) Fictive benchmark study in predictive modeling: an estimated accuracy is obtained for each of the four data sets (rows) and three methods under investigation (columns). Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method in the given row of the table, complicating the aggregation of its performance over all rows required to obtain an overall performance estimation.
Benchmark data set Method 1 Method 2 Method 3
@@
Table 1 depicts a hypothetical summary of the results of your study in each simulation repetition/benchmark data set, denoting the undefined values by a blue “NA”. You may now wonder how to accurately determine the methods’ performances from this incomplete information. Note that Table 1 first of all highlights an important distinction that must be made between simulation and benchmark studies regarding the consequences of method failure. In a simulation study focusing on assessing the methods’ performances in estimating an estimand, a method typically produces a single estimate for the entire data set in the given repetition (i.e., row in Table 1(a)). Its performance is estimated as an aggregate over all rows in Table 1(a). In this setting, method failure first leads to an undefined estimate value which then complicates the assessment of the affected method’s performance. In a benchmark study focusing on the predictive performance of methods, on the other hand, a method makes a prediction for each row of the corresponding test data for a given benchmark data set (based on the decision rule generated on the training data). The method’s performance (e.g., accuracy) is then measured directly in each real data set (i.e., each row in Table 1(b)) and its overall performance is averaged over all rows of Table 1(b). Method failure results directly in an undefined performance value.
Looking at Table 1, it is tempting to view method failure as missing data: in each case, the row summarising results contains “holes” where results were expected. We caution against this. Missing values are data that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023). In our context, although we intended to record a value, method failure means that the value does not exist for that particular data set. This is similar to data truncation by death as opposed to the inability to collect data due to the non-appearance of a patient at the doctor’s office (missing data). As a result, some of the most commonly used approaches to handling method failure such as imputation, being “missing data inspired”, are inadequate in most cases. Furthermore, only few are aware that the typical ways and reportings of discarding data sets associated with failure (either for all or the failing methods only) do not enable meaningful method comparisons. We illustrate this in detail in Sections 3 and 4.
In the subsequent part of this paper, we use the term “undefined” instead of “missing” to emphasize the important distinction between method failure and the “regular” missing data context. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
@@
The rest of this paper is structured as follows. In Section 2, we review and summarize the different manifestations of method failure in published comparison studies. In particular, we discuss how at the core, they often result from a complex interplay of multiple factors, and that an adequate handling requires thorough investigation on a case-by-case basis. Building on this, Section 3 is a compilation of practical recommendations for handling method failure which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure. This approach, if based on thorough and realistic considerations, enables the aggregation of performance across data sets even when method failure occurs. However, considering possible time constraints, we also discuss “slimmed-down” ways in which authors can still implement our recommendations. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
2 Method failure in the literature
In our own experience, method failure occurs frequently when conducting comparison studies in statistics and predictive modeling. As already described, a look at the literature suggests that others also encounter this problem, either because method failure is explicitly mentioned in the text or because signs of method failure can be found in the code provided by the authors for reproducibility. Section 2.1 briefly outlines how method failure manifests and is handled in a selection of published studies from the classical statistics and predictive modeling literature. In Section 2.2, we argue that at the core, method failure is a complex interplay of multiple factors, deserving particular attention. Finally, Section 2.3 briefly categorizes different procedures for handling method failure used in these studies and provides a first outlook on the recommendations provided in Section 3.
2.1 Reporting and handling of method failure from the literature: vignettes
The following comparison studies should be viewed as positive examples, being among the few that report the occurrence of method failure, often alongside reasons and strategies for handling the resulting undefined values when summarizing and interpreting the results. The strategies adopted in these studies inspire—but do not completely overlap with—our recommendations given in Section 3. In these cases, we do not intend to criticize the applied approaches but aim to further improve the handling and reporting of method failure for future work.
van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable” recommendations for logistic regression. They particularly observe that separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates) are commonly omitted from the analysis since they lead to convergence issues of the likelihood maximization process. However, van Smeden et al., (2016) demonstrate that the removal of these data sets, in which the association between the covariates and outcomes is particularly strong, substantially affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small.
In Voeltz, (2024), a simulation study is conducted to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity. Two of the three methods under investigation are occasionally affected by method failure due to non-convergence (while the third one does not employ an iterative process). In the case of non-convergence, one of the two does not return an error but outputs the current, non-converged estimate. As these outputs are often used in practice nevertheless, the authors also include them in the performance comparison. For the remaining method affected by non-convergence, they discard the data sets for which it fails and use the proportion of simulation repetitions in which the methods converge as an additional performance measure.
An extreme example of excluding performance data is to remove an entire method. This is done by Ruxton and Neuhäuser, (2013) in a simulation study to compare the empirical coverage of methods commonly used to calculate confidence intervals for the odds ratio. Given a binary outcome and binary exposure variable, some methods fail due to division by 00 if the corresponding $2\times 2$ contingency table contains one or more zero-entries.
Masaoud and Stryhn, (2010) compare the performance of different marginal and random effects estimation methods for the analysis of binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo) in a simulation study. With data simulated using two different data-generating mechanisms (i.e., stemming from marginal and random effects models), methods are compared regarding bias and coverage of the confidence intervals (as well as efficiency for data simulated from the marginal model). The authors encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010), for which the handling varies for the different methods affected. They either (i) change the software from R to SAS, (ii) switch the optimization technique used internally by the method, or (iii) for GEE, resort to using model-based standard errors if the method fails to produce meaningful robust standard errors. However, if method failure cannot be solved using these measures, they discard the corresponding data set.
Dunias et al., (2024) conduct a simulation study to compare the performance of four hyperparameter tuning procedures (including $k$-fold cross validation and the one-standard error rule for tuning) for prediction methods developed using predictive modeling methods such as Ridge, LASSO, or Random Forest. For some data sets, LASSO does not select any predictor variable, leading to a constant predictor. Resulting in the inability to assess the performance in terms of the calibration slope, the “[missing] calibration slopes [are] replaced by the maximum value of calibration slope for the corresponding scenario” (Dunias et al., , 2024). Note that this corresponds to imputation by the worst “observable” value since higher values of the calibration slope generally indicate that “risk predictions are too narrowly distributed” (Dunias et al., , 2024).
An approach to handling method failure by imputation is applied by Bischl et al., (2013) in their comparison study of $11$ methods for classification on $49$ real and simulated (benchmarking) data sets. They use misclassification error as performance measure combined with $100$ subsampling (also called “Monte Carlo cross-validation”) repetitions. When encountering method failure, they impute based on the proportion of repetitions in which the affected method fails. If below $20\%$, they impute by sampling from a normal distribution that is estimated from the remaining performance values of that model. This way of imputing serves to adhere to the assumptions of the model used to test the generalization performance. If a method fails for more than $20\%$ of repetitions, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013).
Herrmann et al., (2021) perform a benchmark study to compare the predictive performance of survival time prediction methods based on high-dimensional multi-omics data using real cancer data sets. They encounter method failure in several cross-validation folds for several methods (including one method with excessive runtimes). Inspired by Bischl et al., (2013), they impute undefined performance values based on the proportion of “failed” cross-validation folds. If it is above $20\%$, they impute by the “worst possible value” of the corresponding metric, else, they impute by the mean performance of the method in the remaining “successful” folds. Note that the latter is equivalent to discarding those data sets in the performance assessment for which the given method fails. In a subsequent project, Nießl et al., (2022) investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. They consider alternative procedures including imputation by the worst possible value and by a value that weights the method’s overall proportion of undefined performance values and its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
Another example can be found in the work of Hornung et al., (2024). They assess the performance of prediction methods that can handle block-wise multi-omics data by artificially inducing specific patterns of block-wise missing values into initially complete multi-omics data. Three of the considered methods fail for some data sets (with a maximum proportion of $34\%$). In some rare cases, failure is caused by memory issues. More commonly, however, the methods are “not applicable” (Hornung et al., , 2024) if the training and test data sets differ in the induced pattern of missing data blocks in certain ways. The resulting undefined performance values are addressed by discarding the corresponding data sets for all methods with the reasoning that discarding them only for the failing methods would lead to confounding by the induced pattern of block-wise missing values. Note that Hornung et al., (2024) argue that instances of memory issues are likely caused by a specific missingness pattern induced into the training data.
Fernández-Delgado et al., (2014) perform an extensive comparison study of supervised classification methods (“classifiers”) from several model families. The classifiers are assessed in $121$ (real) benchmarking data sets in combination with $4$-fold cross-validation and “accuracy” as performance measure (which is then averaged over all benchmarking data sets). Additionally, they compute the Friedman ranking (Sheskin, , 2011) based on accuracy. They encounter method failure in $2.1\%$ of classifier-data set combinations. As reasons for failure, they list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014). Note that equal predictor values of all observations in the training data lead to the failure of the corresponding method since this violates its requirements. Memory issues also lead to method failure with large data sets, while particularly low numbers of observations in some classes in small data sets lead to errors for some classifiers. The authors discard the corresponding data sets for the failing methods to obtain each classifier’s accuracy. For the Friedman ranking, which requires an identical number of accuracy values for all classifiers, they impute undefined accuracy values with the mean value across all classifiers for this given data set.
Gijsbers et al., (2024) introduce a benchmarking tool for automated machine learning (“AutoML”) frameworks, which are considered as the “methods” under comparison. They conduct a comparison study of nine such methods based on various data sets using the Area Under the Curve (AUC) as performance metric. They categorize the reasons for failure into (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML framework. Categories (i) and (ii) are reported as the most common. Gijsbers et al., (2024) argue against ignoring undefined performance values, stating that they do not occur at random but are correlated with data set characteristics. They impute by the performance of a constant predictor.
2.2 Typical manifestations of method failure
Summarizing from the previous Section 2.1, there are common ways in which method failure often reveals itself to authors of comparison studies; see Table 2 for an additional overview.
Table 2: Common manifestations of method failure
Manifestation Explanation Examples from literature
Non-convergence
and other
calculation issues
Method cannot perform required calculations (within maximum number of iterations), therefore fails to produce a stable output. Cannot be solved by increasing memory or time budgets.
$\rightarrow$ Example: - division by zero - attempt to invert singular matrix $\rightarrow$ Can result in: - error message - non-meaningful output such as NA, NaN, output of $\infty$, or inflated estimation (for estimation task)
van Smeden et al., (2016)
Fernández-Delgado et al., (2014)
Memory
issues
Method consumes more memory than available to perform required computations.
$\rightarrow$ Can result in: system crashes (“fatal error”)
Hornung et al., (2024)
Fernández-Delgado et al., (2014)
Runtime
issues
Method fails to perform required computations within pre-specified or feasible time limits.
$\rightarrow$ “Failure by design” of study
$\rightarrow$ Real users would likely obtain valid output
Herrmann et al., (2021)
Gijsbers et al., (2024)

### v2

# Rethinking the handling of method failure in comparison studies

@@ Abstract
Comparison studies in methodological research are intended to compare methods in an evidence-based manner to help data analysts select a suitable method for their application. To provide trustworthy evidence, they must be carefully designed, implemented, and reported, especially given the many decisions made in planning and running. A common challenge in comparison studies is to handle the “failure” of one or more methods to produce a result for some (real or simulated) data sets, such that their performances cannot be measured in those instances. Despite an increasing emphasis on this topic in recent literature (focusing on non-convergence as a common manifestation), there is little guidance on proper handling and interpretation, and reporting of the chosen approach is often neglected. This paper aims to fill this gap and offers practical guidance on handling method failure in comparison studies. After exploring common handlings across various published comparison studies from classical statistics and predictive modeling, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases. We then recommend a different perspective on method failure—viewing it as the result of a complex interplay of several factors rather than just its manifestation. Building on this, we provide recommendations on more adequate handlings of method failure derived from realistic considerations. In particular, we propose considering fallback strategies that directly reflect the behavior of real-world users. Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two exemplary comparison studies.
@@ 1 Introduction
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure.” Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. In particular, published reviews have shown that method failure goes unreported in the majority of cases. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur(Hinds et al., , 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) report convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility, even when not reported in the manuscript. This is often clear through the use of explicit error-handling mechanisms (e.g., functions try() or tryCatch() in R or try-except constructs in Python).
@@
To illustrate the difficulty of handling method failure adequately, imagine a researcher conducting a comparison study in statistics or predictive modeling to compare the performance of three methods based on four simulation/benchmark data sets. In the context of classical statistics, they likely focus on comparing the methods’ performances (e.g., bias) in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s),” and base the analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, they are likely to compare the methods’ predictive performances (e.g., accuracy), that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmarking data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies,” respectively (though admitting that there may be intermediate situations, e.g., simulation studies investigating predictive performance on artificial data).
@@
During the experiments, the researcher encounters the failure of some methods in one or more of the following ways: (i) they receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) their system crashes, or (iii) the experiment runs for an excessively long time, forcing the researcher to abort.
While the majority of real published comparison studies do not report method failure, there are positive examples from classical statistics and predictive modeling in which the occurrence (and oftentimes also the applied handling) of method failure is described. For instance, van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable (EPV)” recommendations for logistic regression. All three studies featured, exploring the effect of EPV on performance measures such as bias and coverage, encounter separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates), which lead to convergence issues of the maximum likelihood process. That is, convergence is not achieved within the pre-specified number of iterations, or the process converges to a point that is not the maximum likelihood estimate (van Smeden et al., , 2016). In a study from the statistical learning field, Gijsbers et al., (2024) conduct a comparison study of nine automated machine learning methods (“AutoML frameworks”). They encounter and report method failure in the forms of (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML method (Gijsbers et al., , 2024).
Table 1: Overview table of (a) a fictive simulation study across four repetitions (rows) and (b) a fictive benchmark study across four benchmark data sets (rows) for three methods (columns). Method failure results in undefined values for the affected combination of data set and method, marked by “NA.”
(a) Fictive simulation study: overview of estimates, with the true value being equal to $4$. Method failure first leads to an undefined estimates value for the given method–data set combination, complicating the derivation of bias for that method.
@@
(b) Fictive benchmark study: overview of estimated accuracies across all method–data set combinations. Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method–data set combination.
Benchm. data set Method 1 Method 2 Method 3
@@
Regardless of the form in which a researcher encounters method failure, it leads to NA for the associated method-data set combination where otherwise an estimate (for simulation studies) or a performance value (for benchmark studies) is expected. Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetition/benchmark data sets and methods. In simulation studies, the NAs complicate the derivation of performance (e.g., bias as the mean deviation of the estimates from the true value across all repetitions). In benchmark studies, on the other hand, they hinder the aggregation of performance values into an average performance (e.g., average estimated accuracy as the mean fraction of correct predictions across all benchmark data sets). In both cases, method failure complicates the assessment and especially the comparison of performance of the methods under investigation such that the researcher finds themself wondering how to proceed adequately.
The above-shown examples of published comparison studies alone already illustrate different viewpoints on method failure and correspondingly different ways in which it is handled in practice. van Smeden et al., (2016) observe that separated data sets are commonly removed from the analysis. This, however, affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small (van Smeden et al., , 2016). Gijsbers et al., (2024) argue against discarding data by describing that method failure is typically correlated with data set characteristics. Instead, they impute by the performance of a constant predictor.
Examining additional published studies, which we do throughout this paper, shows that imputation and discarding data sets associated with failure are among the most common practices. However, it also becomes clear that they are only two of many handlings applied in practice and that others, such as modifying the affected methods’ implementations and changing software, are common as well—an indicator that proper guidance is commonly lacking. Our paper aims to fill this gap and provides detailed instructions on how to handle method failure. In particular, we illustrate why usual handlings of method failure derived from treating the resulting NAs as regular “missing data,” and especially imputation and discarding data, are inappropriate in most settings (leading us to refer to “undefined” rather than “missing” performance values in instances of method failure). As an alternative, we promote considering method failure from a different viewpoint, namely as the result of a complex and individual interplay of multiple factors—which we use to derive alternative and more appropriate measures that can be applied when encountering method failure.
Our paper targets both the planning and the execution stage of a comparison study. While many instances of method failure can only be directly observed—and therefore investigated—after the study is executed, certain considerations should be made while planning to anticipate and ideally reduce the potential for failure.
@@
The rest of this paper is structured as follows. In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values are usually inadequate when encountering method failure. Building on this, Section 3 is a compilation of alternative recommendations for handling method failure, which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs and on top of that, reflects the behavior of real method users across many settings. Considering possible time constraints, we also discuss “slimmed-down” ways in which authors can implement part of our recommendations if unable to follow them completely. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
2 Method failure in published comparison studies
In Section 2.1, we summarize common manifestations and handlings of method failure reported in the literature. Since some authors focus on how method failure is handled rather than explaining how it occurred, the corresponding studies are discussed only in Section 2.1.2. An overview of all study contexts, manifestations, and handlings of method failure are provided in Table 2.
2.1 Status quo: occurrence and handling of method failure
Comparison study     Manifestations     Handlings
        Resolve Summarize despite
        method failure method failure
Author Context of the study     non-convergence (and others) runtime issues
memory issues
Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Performance measure
van Smeden et al., (2016)a Simulation study to investigate why different published simulation studies offer conflicting minimal “events per variable (EPV)” recommendations for logistic regression. Features three published simulation studies exploring the effect of EPV on performance measures.     $\times$     $\times$
Gijsbers et al., (2024) Comparison study of nine automated machine learning (“AutoML”) methods.     $\times$ $\times$
$\times$
$\times$
Zapf et al., (2024) Simulation study to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity.     $\times$     $\times$ $\times$
Ruxton and Neuhäuser, (2013) Simulation study on the empirical coverage of eight methods for calculating confidence intervals for the odds ratio for binary outcome and exposure.     $\times$     $\times$
Hornung et al., (2024) Benchmark study using thirteen multi-omics data sets to assess the performance of seven prediction methods for multi-omics data that can handle block-wise missingness by artificially inducing specific patterns of block-wise missing values into initially complete data.     $\times$
$\times$
$\times$
Masaoud and Stryhn, (2010) Two simulation studies to compare the performance of seven marginal and random effects estimation methods for binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo).     $\times$     $\times$ $\times$ $\times$
Crowther et al., (2014) Simulation study to assess the performance of a multilevel mixed effects parametric survival model using adaptive and nonadaptive Gauss-Hermite quadrature.     $\times$     $\times$
Dunias et al., (2024) Simulation study to compare the performance of five hyperparameter tuning procedures for clinical prediction models built upon statistical learning methods.     $\times$     $\times$
Fernández-Delgado et al., (2014) Benchmark study of 179 supervised classification methods across 121 benchmark data sets.     $\times$
$\times$
$\times$ $\times$
Herrmann et al., (2021) Benchmark study on the predictive performance of thirteen survival time prediction methods based on eighteen high-dimensional multi-omics data.     * $\times$ *     $\times$ $\times$
Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$ $\times$
Nießl et al., (2022)b Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. Considered alternative procedures include imputation by the worst possible value and by a value that weighs the method’s overall proportion of method failure against its performance in the remaining folds.     * * *     $\times$
  • a
    van Smeden et al van Smeden et al., (2016) observe that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
  • b
    Nießl et alNießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et alHerrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
    The authors did not report details on the manifestation of method failure.
Table 2: Selection of comparison studies from statistics and predictive modeling that report the occurrence and handling of method failure.
2.1.1 Common manifestations of method failure